### v1

# NaP2: A Benchmark for Naturalness and Privacy-Preserving Text Rewriting by Learning from Human

Abstract
Increasing concerns about privacy leakage issues in academia and industry arise when employing NLP models from third-party providers to process sensitive texts. To protect privacy before sending sensitive data to those models, we suggest sanitizing sensitive text using two common strategies used by humans: i) deleting sensitive expressions, and ii) obscuring sensitive details by abstracting them. To explore the issues and develop a tool for text rewriting, we curate the first corpus, coined NaP2, through both crowdsourcing and the use of large language models (LLMs). Compared to the prior works based on differential privacy, which lead to a sharp drop in information utility and unnatural texts, the human-inspired approaches result in more natural rewrites and offer an improved balance between privacy protection and data utility, as demonstrated by our extensive experiments.
1 Introduction
Data sharing and information dissemination between AI models are pivotal in the AI era, particularly since the emergence of Large language models (LLMs). The remarkable performance of LLMs benefit from a large amount of shared and publicly available data. However, it is still challenging to balance between data privacy and information utility when training and utilizing such LLMs Pan et al. (2020). Users or downstream applications often interact with commercial LLMs by directly inputting raw text. Such interactions can inadvertently expose sensitive data, such as personally identifiable information (PII), to untrusted service or LLM providers Utpala et al. (2023).
Redaction and anonymization techniques are widely applied to remove PII from texts, but they suffer from three major drawbacks Sánchez et al. (2014). First, after anonymization, mentions of PII are either redacted or replaced by their entity types so that processed texts become unnatural. Downstream applications need to be adapted or fine-tuned to cope with such unnatural texts. Second, it is still possible to recover private attributes from PII scrubbed text via reasoning Mireshghallah et al. (2023); Staab et al. (2023). Third, the presence of blacked-out parts or entity types may raise the awareness of a document’s sensitivity in front of potential attackers.
ORI: I am Cindy .
I am recovering from
ulnar nerve surgery .
PER: I just had surgery.
Human Rewrite:
DEL: I am Cindy.
OBS: Hi, I am Cindy, I am just
recovered from medical treatment
T5 trained on NaP2:
Output: I am recovering from llness.
DP method
DPNR: YYYYYYYYY.
DP-Forward: My name is Cindy and I am
recovering from ulnar nerve surgery.
Table 1: An example of rewriting a text (ORI) using deleting (DEL) and obscuring (OBS) as the strategies based on personal information (PER).
Alternatively, differential privacy (DP) provides a theoretical privacy guarantee for data release or dissemination mechanisms Dwork (2006). Prior works sanitize texts by perturbing texts either at the word-level or the sentence-level Mattern et al. (2022); Igamberdiev and Habernal (2023); Igamberdiev et al. (2022a). In order to reach a high-level of privacy guarantee, substantial noise need to injected into texts or their representations so that information utility drops sharply and the meanings of texts are changed significantly (see Table 1). Therefore, determining an optimal trade-off between privacy and utility for data release remains an unresolved challenge.
To address the shortcoming of prior methods, we propose an approach that adopts human text editing strategy inspired by  Strengers et al. (2020), specifically deleting and obscuring to improve the naturalness and utility of rewritten texts while ensuring privacy. As shown in Table 1, given an utterance involving personal information stated in a persona, the strategy deleting simply removes all words mentioning sensitive information from the utterance, while obscuring substitutes sensitive expressions for more abstract and general expressions. Both strategies aim to make rewritten texts as natural as possible such that i) they do not raise the awareness of potential attackers that rewrites are sanitized; and ii) downstream applications can directly process such natural rewrites without fine-tuning their models for any unnatural parts of texts.
To evaluate strategy-specific rewriting models, we construct the first Naturalness and Privacy Preserving Rewriting corpus, coined NaP2, based on the open-domain dialogue corpus PERSONA-CHAT Zhang et al. (2018). We recruit university students to manually rewrite 895 utterances involving personal information as the manual evaluation set. To promote the development of diverse open-source solutions for this task, we apply GPT4 to generate 3900 synthetic examples as the synthetic training set because GPT4 demonstrates the best performance on PERSONA-CHAT among all evaluated models. We also design multiple automatic and human evaluation metrics for this task, including a novel privacy metric Privacy_NLI. It utilizes a Natural Language Inference (NLI) model Liu et al. (2019) to determine if a rewrite entails a personal information or not. The extensive comparative studies between the models trained on our corpus and the state-of-the-art (SOTA) text sanitization methods demonstrate the underlying challenges and yield the following key findings:
  • The T5-Base model Raffel et al. (2020) trained on our corpus is able to achieve a fairly high privacy preservation indicated by a Privacy_NLI of 93.81%. Its performance is even significantly superior than GPT4 according to human evaluation using deleting. In contrast, the competitive DP methods have a Privacy_NLI score lower than 62.14%.
  • The privacy metric Privacy_NLI aligns well with the human judgements by having a Spearman’s ranking correlation of 0.70.
  • GPT4 generates synthetic rewrites with decent trade-off between privacy and utility based on human evaluation, better than GPT-3.5 turbo and the evaluated open-source LLMs in the zero-shot setting. Incorporation of such synthetic data improves the T5-Base model trained on human curated data by 7% in terms of privacy preservation.
2 Preliminary
As our task is closely related to local differential privacy (LDP) Xiong et al. (2020), this section mainly introduces the preliminary concepts of LDP and Context-Aware LDP Acharya et al. (2020).
LDP.
For private statistical data publication, LDP provides strong and provable privacy preservation without assuming that data collectors can be trusted. In this setting, each participant locally perturbs her/his private data with a randomized mechanism and transfers the perturbed output to a data collector. The data collector acquires statistical information from the perturbed data received from all participants without compromising the individual’s privacy.
Definition 2.1 (Pure Local Differential Privacy Kasiviswanathan et al. (2011)).
Let $\epsilon\geq 0$, a randomized algorithm $Q:\mathcal{X}\rightarrow\mathcal{Y}$ is $\epsilon-$locally differential privacy, if for all $x,x^{\prime}\in\mathcal{X}$ and $y\in\mathcal{Y}$,
$\frac{Q(y|x)}{Q(y|x^{\prime})}<e^{\epsilon}$
where $Q(y|x)$ can also be viewed as a conditional distribution. If the privacy budget $\epsilon$ is small or zero, it is difficult or infeasible to distinguish between $x$ and $x^{\prime}$ based on the outputs of $Q$. However, it is questionable if such perturbed data is still useful for data analysts or downstream applications (Mattern et al., 2022).
This definition assumes that all elements in $x$ are equally sensitive and all $x$ share the same privacy constraint $e^{\epsilon}$ regardless of how different $x$ and $x^{\prime}$ are. For NLP applications, Mattern et al. (2022) observe that a tight universal privacy budget leads to substantial grammatical errors produced by word-level DP mechanisms, while a high budget easily compromises individual privacy. Therefore, the privacy budget of $x$ and $x^{\prime}$ should vary depending on the semantic differences between $x$ and $x^{\prime}$ or whether $x$ or $x^{\prime}$ contains sensitive information.
To address the limitations, Acharya et al. (2020) propose context-aware LDP that employs a different privacy constraint based on differences between $x$ and $x^{\prime}$.
Definition 2.2 (Context-Aware LDP Acharya et al. (2020)).
Let ${\mathbf{E}}\in\mathbb{R}^{m\times m}$ be a matrix of non-negative entries and $\epsilon_{x,x^{\prime}}$ denote the $(x,x^{\prime})$th entry of ${\mathbf{E}}$, a random algorithm $Q$ is ${\mathbf{E}}-$LDP, if for all $x,x^{\prime}\in\mathcal{X}$ and $y\in\mathcal{Y}$,
$\frac{Q(y|x)}{Q(y|x^{\prime})}<e^{\epsilon_{x,x^{\prime}}}$
The matrix ${\mathbf{E}}$ can be constructed by using different functions. Metric-based LDP Alvim et al. (2018) can also be viewed as a special case of context-aware LDP by requiring $\epsilon_{x,x^{\prime}}=\epsilon d(x,x^{\prime})$, where $d(x,x^{\prime})$ is the metric between $x$ and $x^{\prime}$.
3 Naturalness and Privacy-Preserving Rewriting
3.1 Problem Definition
Task.
Given an utterance ${\bm{x}}$ and a sentence ${\bm{p}}$ describing personal information, the task of naturalness and privacy-preserving rewriting aims to map ${\bm{x}}$ into a natural sentence ${\bm{y}}$ such that ${\bm{y}}\in\mathcal{Y}^{n}$ does not reveal the personal information in ${\bm{p}}$ and maximally preserves the non-private content in ${\bm{x}}$. We define a natural sentence as one that is grammatically correct, fluent, and does not contain any artifacts such as blacked-out words or special symbols indicating omitted sensitive information. The rewrite space $\mathcal{Y}^{n}$ contains only natural sentences. Compared with differential privacy (DP) mechanisms that prevent privacy leakage during model training Abadi et al. (2016a), this task focuses on privacy-preserving data publishing or privacy protection at inference time.
When sanitizing texts, humans often hide sensitive information by avoiding sensitive words or replacing them with more general or abstract expressions Strengers et al. (2020). We expect machines to adopt similar strategies:
  • Deleting: removing words or phrases in ${\bm{x}}$ that leak personal information specified in ${\bm{p}}$;
  • Obscuring: replacing sensitive words or phrases in ${\bm{x}}$ with more general or abstract expressions to avoid compromising privacy.
Relation to LDP. A probabilistic rewriting model can be viewed as a random mechanism $Q(y|x)$ that maps an input text $x$ into a word sequence $y$ inside a constraint output space $\mathcal{Y}$, which only contains natural texts. Given a pair of semantically similar texts $x$ and $x^{\prime}$, where only $x$ contains sensitive information, a rewriting model $Q(y|x)$ implements metric-based LDP by enforcing the distribution divergence in log-scale between $Q(y|x)$ and $Q(y^{\prime}|x^{\prime})$ to be smaller than $\epsilon d(x,x^{\prime})$. As a result, perturbed texts are more similar than their original counterparts because mentions of private information are either removed or obscured.
Corpus Overview. Our corpus NaP2 consists of a small manually curated dataset for both training and testing (Sec. 3.2), and a large synthetic dataset distilled from GPT-3.5 turbo and GPT4 for training data augmentation (Sec. 3.3). According to our evaluation stated below, human writes with obscuring achieve the best trade-off between privacy and utility, and the naturalness of GPT4 generated texts is on par with that of human rewrites.
3.2 Manually Curated Corpus
The corpus PERSONA-CHAT associates each multi-turn chit-chat with two personas, each of which is a set of sentences describing the corresponding personality. Hence, it is straightforward to measure if an utterance leaks personal information in the relevant persona. From another point of view, a persona can be regarded as a user-specific privacy profile, which states what information needs to be protected. For instance, one user might consider their marital status as sensitive information requiring privacy protection, while another user may not prioritize it.
The manual created evaluation set extends the test set of PERSONA-CHAT with human-authored rewrites. As not all utterances reveal private information in personas, we apply the automatic alignment methods to pair an utterance involving personal information with the corresponding sentence in a persona.
Formally, given a dialogue $\mathcal{D}$, suppose there are $m$ utterances $\mathcal{X}_{i}=\{{\bm{x}}_{1},{\bm{x}}_{2},...,{\bm{x}}_{m}\}$ associated with a persona $\mathcal{P}_{i}=\{{\bm{p}}_{1},{\bm{p}}_{2},...,{\bm{p}}_{n}\}\}$, we aim to compute an alignment score $s_{ij}$ between ${\bm{x}}_{i}\in\mathcal{X}_{i}$ and ${\bm{p}}_{j}\in\mathcal{P}_{i}$ indicating to what degree ${\bm{x}}_{i}$ leaks personal information in ${\bm{p}}_{j}$.
We formulate the computation of alignment scores as an NLI problem. Namely, if ${\bm{x}}_{i}$ entails ${\bm{p}}_{j}$, it is highly likely that ${\bm{x}}_{i}$ leaks information in ${\bm{p}}_{j}$. Specifically, we reuse the RoBERTa model trained on Multi-Genre Natural Language Inference (MNLI) corpus Williams et al. (2018), which is available from Huggingface, to compute the probability of $p(y=\text{entail}|{\bm{x}}_{i},{\bm{p}}_{j})$ as $s_{ij}$. We find out that this simple approach significantly outperforms Sparse-Max and Sharp-Max proposed in Xu et al. (2020) on a random sample of 200 ground-truth pairs. We manually check the candidates among the pairs with a score higher than a threshold and keep only the well aligned ones.
For each selected sentence-persona pair, we recruit annotators from Amazon Mechanical Turk (AMT) to rewrite utterances w.r.t. the aligned persona sentences using both Deleting and Obscuring.
In our preliminary experiments, we observe that even though annotators endeavor to generate decent rewrites, many of them could not clearly identify and strictly stick to the required strategies. Therefore, we prepare a small sample of pairs as a pre-test to select qualified annotators. In addition, we employ a rigorous procedure for quality check. In particular, we wrap up 15 sentence-persona pairs as a batch and ask annotators to rewrite them using the required strategies. Then, we manually check the rewritten batches, we only accept those that are written using the required strategy. The averaged acceptance rate of the rewrites is 47.97%, demonstrating the challenge of collecting a high-quality rewriting dataset with specific rewriting requirements. As a result, we collect 895 pairs annotated with one rewrite per strategy. We further split this corpus into a cross-validation (CV) set, a validation and a hold-out test set with 655, 140 and 100 instances, respectively.
Data Statistics.
We analyze the manually curated corpus using averaged word length in sentences (Len.) and distinct unigrams divided by the total number of words (Dist.) Li et al. (2016). The statistics of the dataset is given in Table 2. Deleting tends to produce more concise rewrites, while obscuring is slightly longer than ORIGINAL sentences. Although the average length increases, the diversity score for obscuring is still ascending, compared with original sentences. This shows the high diversity of word usage using obscuring.
CV Valid Test
Len. Dist. Len. Dist. Len. Dist.
ORI 13.7 0.148 13.6 0.257 13.5 0.248
DEL 8.0 0.190 8.4 0.298 8.5 0.279
OBS 14.1 0.160 13.9 0.266 14.3 0.250
Table 2: Statistics of original sentence (ORI), rewrites with deleting (DEL) and obscuring (OBS) on the CV set, validation and test set of the manually curated dataset, using average length (Len.) and distinct token (Dist.)
3.3 Synthetic Data Augmentation
We employ the RoBERTa NLI model to align utterances with persona sentences in the training set of PERSONA-CHAT and keep only the pairs with an entailment probability above 0.3. This threshold leads to high recall low precision alignments so that GPT4 is employed to check if there is indeed a privacy leakage. Among them, we randomly sample 3900 pairs to generate synthetic rewrites by using GPT4. The resulting dataset is used to augment the training set of the manually created corpus to mitigate the data scarcity issue.
Prior studies show that GPT4 is one of the strongest few-shot learner Brown et al. (2020). Therefore, we carefully design prompts and in-context examples to use it for privacy-aware rewriting. Given an utterance-persona pair, we use the following prompt for a selected rewriting strategy.
Rewrite this sentence, $<$deleting / obscuring$>$ any private information. Example rewrites are: <\$absentcurrency-dollar<\\$< \$$IN-CONTEXT\_EXAMPLES$$>$ Only return the rewritten sentence, nothing else. Private information present is: [\$PERSONA]. Sentence to rewrite is: [\$UTTERANCE].
where \$X denotes a placeholder for the corresponding information. The $k$ in-context examples are selected from a combination of the validation set of the manually curated corpus and a set of non-sensitive utterances which do not leak personal information. Each of the in-context examples in the validation set contains an utterance, a persona sentence, and a human rewrite using the given strategy, while an example from the non-sensitive set includes only an utterance. The in-context examples are found by $k$-nearest neighbour search using the sentence embeddings of utterances Reimers and Gurevych (2019). In this work, given an utterance, we select the top-$1$ most similar example from the validation set and one example from the non-sensitive set. The latter is used to instruct GPT4 that it should not rewrite an utterance if there is no privacy leakage detected.
3.4 Human Evaluation
Three university students are recruited to check their quality on a set of 100 instances sampled from the test set of the manual corpus. Hence, an utterance-persona pair in the sample includes a human rewrite, a rewrite from GPT-3.5 turbo and GPT4 respectively. For each rewrite, a student is instructed to answer the following questions from the perspectives of privacy leakage (Q1), semantic relevance (Q2) and naturalness (Q3) which is detailed in Appendix A.1.
Each question is answered by three university students. To deal with possible disagreements, we take the majority vote as the final answer.
In order to use a score to summarize the performance w.r.t. each criteria, we calculate the percentage of choosing the option (a) as the majority vote for each question above on the human evaluation test set, referred to as SPrivacy, SRel, and SNatural. They indicate the percentage of rewrites having no privacy leakage, complete semantic relevance, full naturalness, respectively.
SPrivacy SRel SNatural
Human_deleting 82.00% 76.00% 95.00%
GPT3.5_deleting 34.00% 94.00% 72.00%
GPT4_deleting 49.00% 92.00% 99.00%
Human_obscuring 81.00% 97.00% 98.00%
GPT3.5_obscuring 61.00% 90.00% 95.00%
GPT4_obscuring 66.00% 95.00% 99.00%
Table 3: Comparison between GPT-3.5 turbo, GPT4, and human rewrites.
Figure 1: Human evaluation of privacy leakage.
To understand the quality of rewrites in our corpus, we compare GPT4 outputs with those of GPT-3.5 turbo using the same prompts, as well as with human rewrites. The key results are summarized in Table 3. Human rewrites achieve the highest level of privacy protection with both strategies, outperform the best rewriting model GPT4 by at least 15%. Human rewrites with obscuring achieve the best balance between privacy and utility in comparison with alternative methods. Both OpenAI models completely preserve personal information in over 60% of utterances by using obscuring, but struggle to implement the deleting strategy for the same purpose. A close investigation on the percentages of individual Q1 answer in Fig. 1 demonstrates that both models fail to delete private expressions completely in over 34% of the utterances involving sensitive information. GPT-3.5 turbo is significantly worse than GPT4 in terms of sanitization. Only a small proportion of the errors are attributed to applying an incorrect strategy.
4 Experiments
4.1 Rewriting Models
We compare the SOTA privacy-preserving rewriting models DPNR Lyu et al. (2020), DP-Forward Du et al. (2023), and the zero-shot LLMs with the T5-Base models fine-tuned on our corpus, with or without synthetic data augmentation. The word-level DP method DPNR and the sentence-level DP model DP-Forward are fine-tuned based on a T5-Base model, which is pre-trained to map inputs to their outputs. All implementation details can be found in Appendix A.3. DPNR. It stands for Differentially Private Neural Representation, which applies Laplace noise to distributed representations of words in order to randomly drop sensitive words or replace sensitive words with non-sensitive ones. We compare the cosine similarity between each word in an input utterance with those in the corresponding persona, and pick the top-$k$ most similar ones.
DP-Forward. This method perturbs embedding matrices and multi-head attention layers during each forward pass of a language models by achieving a sentence level LDP. When adapting this approach to T5-Base for inference, we mainly perturb embedding matrices, because the DP mechanism for attention layers is mostly useful for protecting privacy at the training time.
Method Privacy_NLI SPrivacy ROUGE-1 ROUGE-Lsum
DPNR 62.14% 25.00% 92.79% 92.79%
DP-Forward 36.42% 0.00% 99.91% 99.91%
DP-Prompt 62.86% 0.00% 42.18% 41.89%
DP-BART 78.22% 1.00% 44.01% 43.15%
FLAIR-SCRUBBING 56.43% 0.00% 67.75% 67.89%
T5_zeroshot-deleting 70.0% 10.00% 16.62% 12.61%
T5_zeroshot-obscuring 45.00% 45.00% 29.58% 23.80%
Llama-13B_zeroshot-obscuring 79.28% 16.00% 40.86% 40.12%
Llama-13B_zeroshot-deleting 77.14% 14.00% 68.28% 67.53%
Llama-paraph-obscuring 82.86% 31.00% 21.72% 20.05%
Llama-paraph-deleting 76.42% 16.00% 56.29% 54.91%
GPT-3.5-obscuring 87.14% 61.00% 66.66% 65.76%
GPT-3.5-deleting 74.29% 34.00% 69.13% 68.48%
GPT-4-obscuring 92.14% 66.00% 73.24% 72.63%
GPT-4-deleting 90.0% 49.00% 77.48% 77.08%
T5-NaP2-GPT4 93.81% 72.00% 73.01% 72.78%
Table 4: Evaluation and comparison of baseline methods.
Llama-paraph. Mattern et al. (2022) points out the limitations of word-level LDP and propose to paraphrase input texts with lower temperature to achieve a sentence-level LDP. We implement this approach by using Llama-13B.
DP-Prompt.Utpala et al. (2023) utilizes zero-shot prompting and large language model to generate document paraphrasing to prevent author de-anonymization attack which comprise the privacy of text owner.
DP-BART. The method is a privatized text rewriting system incorporates LDP. The system leverages the LPD paradgram to perform model rewriting using BART model to protect input data which tackles same challenge like us.
FLAIR-SCRUBBING. we also adapt the scrubbing method used in  Lukas et al. (2023) as our baseline. We employ FLAIR-SCRUBBING as our method to test if this automatics method can effectively remove private information from sentence.
Zero-Shot LLMs. To compare with the LLMs fine-tuned on our corpus, we apply the same prompts to the same pre-trained LLMs without any training. Specifically, we consider T5-Base, Llama-13B, GPT-3.5 turbo and GPT4 and apply the prompt template introduced in Sec. 3.3. To distinguish from the fine-tuned models, the T5-Base and Llama-13B in the zero-shot setting is referred to as T5_zeroshot and Llama-13B_zeroshot, respectively.
T5-NaP2. By using the same prompts as the zero-shot version, we fine tune T5-Base on the training set of the manually curated corpus, with or without augmenting them with synthetic data. The prompts are similar to those used by zero-shot models detailed in A.2.
T5-NaP2-DP. To simulate the use cases that the training data of the rewriting models contains sensitive information, we apply DP-SGD Abadi et al. (2016a) when fine-tuning the T5-Base model in order to understand to what degree the DP mechanism impacts the inference quality of the rewriting models and shed light on future research directions.
4.2 Evaluation Details
Prior studies focus on protect data privacy from membership inference attacks, reconstruction attacks, and sensitive attribute attacks etc. Mattern et al. (2022). However, almost all of them focus on privacy preservation at the training time. In contrast, our target task is concerned with i) if a rewrite reveals personal information in a given persona, ii) preservation of non-sensitive content, and iii) naturalness of rewrites. Compared with the prior studies based on DP mechanisms, our setting is more close to that of natural language generation (NLG) tasks. Therefore, we evaluate the outcomes of the rewriting models by using NLG motivated automatic and human evaluation.
For human evaluation, we use the same questionnaires and the metrics introduced in Sec. 3.4 and ask annotators to answer each question in order to obtain the majority votes.
For all experiments involving model fine-tuning, we conduct five folds cross validation (CV) on the CV set of the manually curated corpus. In order to understand the usefulness of synthetic data, we also conduct experiments with the same models that augment the training set in each fold with 3,900 synthetic instances generated by GPT4.
4.2.1 Automatic Evaluation Metrics.
Privacy Leakage. We propose a novel metric, called Privacy_NLI, by using the RoBERTa model trained on the MNLI corpus, to infer to what degree it is possible to infer personal information in personas. As the NLI model classifies a pair of input texts into entailed, contradicted, or neutral, we adopt $P(\text{entailed}|{\bm{x}},{\bm{p}})$ as the score of $privacy\_leakage$, e. Hence, we consider Privacy_NLI as 1- $privacy\_leakage$, denoting the privacy preserved by our method. The higher the metric, the more private information is preserved.
Semantic Relevance. For assessing the preservation of semantic content, we consider ROUGE-1 and ROUGE-Lsum Lin (2004) to compare generated rewrites with the corresponding references.
4.3 Results and Discussions
SPrivacy SRel SNatural
Human_deleting 82.00% 76.00% 95.00%
Llama-13B _deleting 54.00% 49.00% 87.00%
T5-NaP2-GPT4 _deleting 72.00% 91.00% 95.00 %
DPNR 1.00% 0.00% 19.00%
Human_obscuring 81.00% 97.00% 98.00%
DP-Prompt 0.00% 1.00 % 0.00%
DP-BART 1.00% 10.00% 2.00%
FLAIR-SCRUBBING 0.00% 1.00% 0.00%
Llama-13B _obscuring 12.00% 14.00% 86.00%
T5-NaP2-GPT4 _obscuring 53.00% 93.00% 98.00%
Table 5: Human evaluation of the SOTA models.
Efficacy of NaP2. Table 4 reports the evaluation of all methods. T5-Base fine tuned on the human rewrites and the synthetic data using both strategies outperform the DP based methods and zero-shot LLMs by a wide margin. DPNR preserves more privacy than DP-Forward, but results in a dramatic drop of information utility. The generated texts often have completely different meanings and have substantial grammatical errors, though some of them are still fluent. In contrast, DP-Forward mostly copies inputs to outputs but rarely hide sensitive information. Llama-paraph produces frequently irrelevant texts, hence have fairly low ROUGE-1 and ROUGE-Lsum scores. Besides, for convention personally identifiable information scrubbing method FLAIR-SCRUBBING, it can not effectively remove the private information in open-ended domain, only 40.71% examples are successfully removing PII tokens. For DP-Prompt and DP-BART, even Privacy_NLI are outperformed than other baseline models, the paraphrasing impairs the semantic of original sentence leading to low ROUGE-1 score.
We further investigate the rewriting quality w.r.t. each strategy based on human evaluation. We use the T5-Base model trained on the human rewrites and the synthetic data with both strategies, and apply it on the hold-out test set of each strategy. Table 5 shows that the T5-Base model achieves superior performance over the baselines with both strategies. The naturalness of all generated rewrites is on par with that of human rewrites. Both zero-shot Llama-13B models perform better than the best DP method DPNR, which mostly perturbs non-sensitive contents or yields repeated words. The overall results are encouraging for a wide range of applications on edge devices, because our corpus is not huge and T5-Base contains only a few million parameters, which is a few hundred times smaller than Llama-13B, GPT-3.5 turbo and GPT4.
Alignments between Automatic metrics and Human Evaluation. We compare the ranking using Privacy_NLI with the corresponding human judgements in Table 4. T5-NaP2-GPT4 obtains the highest 1-Privacy_NLI of 93.81% in automatic evaluation, matching the highest SPrivacy with 72.00%. The results are aligned well among the rewriting models using the obscuring strategy. However, Privacy_NLI does not rank all rewriting models using deleting in the same manner as humans. To quantify the alignments, we calculate a Spearman’s ranking correlation of 0.70 between Privacy_NLI and SPrivacy among all models to show the effectiveness of Privacy_NLI. The correlation between the models using obscuring reaches even 0.83.
Usefulness of the Synthetic Data. Table 6 shows the result of using synthetic data for training rewriting models. We compare two different strategies: deleting and obscuring. The results shows that the model performs better with the synthetic data for both tasks. In particular, the model preserves more non-personal information compared to human rewrites in the deleting task. With the synthetic data for training the models, the model performance is 7% better than the non-synthetic data model in terms of deleting. The biggest gain of the synthetic data is obtained for improving the privacy protection of the rewriting model using obscuring.
SPrivacy SRel SNatural
Human_deleting 82.00% 76.00% 95.00%
T5-NaP2-GPT4 _deleting 72.00% 91.00% 95.00%
non-Syn_deleting 65.00% 92.00% 93.00%
Human_obscuring 81.00% 97.00% 98.00%
T5-NaP2-GPT4 _obscuring 53.00% 93.00% 98.00%
non-Syn_obscuring 4.00% 92.00% 93.00%
Table 6: Human evaluation results with and without synthetic data.
5 Related Work
The field of controllable text style transfer focuses on modifying specific attributes in texts, such as formality Briakou et al. (2021) and sentiment Li et al. (2018a, 2022) while preserving the core semantic content. The advancement of text rewriting tasks is heavily dependent on the availability of high-quality corpora to assess generation quality. For example,  Rao and Tetreault (2018) collected a large-scale corpus GYAFC for initiating the research of formality style transfer to rewrite formal language. As for our task sensitive to privacy, which demands sophisticated alignment in rewriting utterances, the construction of a specialized corpus for high-quality privacy-sensitive rewrites are crucial.
There is a growing interest in protecting user privacy Chen et al. (2020); Tigunova et al. (2019); Xu et al. (2019); Bevendorff et al. (2019) in NLP tasks. One way of protecting privacy is to implicitly remove the information in decision models, for example perturbing the representations via adversarial training  Li et al. (2018b); Elazar and Goldberg (2018); Barrett et al. (2019) or differential privacy Fernandes et al. (2019); Bo et al. (2019). In text rewriting which is close to our rewriting approach, local differential privacy are recently adapted to protect the data by adding customized noise Igamberdiev et al. (2022b); Igamberdiev and Habernal (2023). Such adaptations in rewriting system mitigate the privacy leakage risk of original input however result in complete semantic change of inputs as the noise is independently drawn from the data and task. We consider a more generalised rewriting setting where the naturalness and general meaning of sentence are preserved.
Another series of work suggested to generate new sentences with less sensitive information Emmery et al. (2018); Xu et al. (2019). Following this approaches, the setting of our work is more general since we use open-domain sensitive personal information as the control signal for rewriting. Moreover, our corpus is flexible in the way that it supports two strategies for rewriting, which is of the interest for the style transfer research community Strengers et al. (2020).
6 Conclusion
We introduce the task of naturalness and privacy-preserving text rewriting and collect a corpus NaP2 based on PERSONA-CHAT. The fundamental concept involves training models to learn human strategies, namely deleting and obscuring, for inference-time privacy. The T5-Base model trained on our corpus outperforms competitive zero-shot LLMs and DP methods by a wide margin. This work paves the way for future research on LLM-based rewriting techniques with a new focus on naturalness preservation.
Ethical Statement
In this paper, we align our research practices with the principles outlined in the ACL Code of Ethics, fully endorsing its values. Our investigation has been conducted in compliance with these ethical standards.
The creation and assessment of NaP2 have been conducted with a keen awareness of ethical considerations, especially regarding the involvement of human annotators. The necessity for human-annotated data to train conditional independence classifiers in our method is recognized as demanding significant effort. We have taken careful measures to ensure that this process is ethically sound, honoring the annotators’ contributions by respecting their time and providing equitable compensation.
Moreover, the central objective of NaP2 is to assess the relevance of generated responses in relation to their persona information and the difference between human evaluation and proposed automated metrics. The system is engineered to assign scores on a continuous scale from 0 to 1, with higher scores denoting greater relevance. It is designed to yield only these scores, without generating any information that could be deemed harmful or violate privacy.
Limitation
Due to budgetary constraints associated with this project, we were unable to engage a vast number of annotators to rewrite the extensive dialogue datasets with respective rewrite strategies. Consequently, NaP2 we compiled is somewhat limited in scope. While NaP2 possesses sufficient volume to validate the core assertions of our study, it might not fulfill the expansive needs of commercial deployments. Industrial entities interested in utilizing our dataset could potentially address this limitation by adopting prompt tuning techniques or employing additional annotators to expand the dataset in accordance with our outlined methodology.
Our evaluation metric is specifically designed to assess the relevance of the generated responses. Although it demonstrates superior performance over baseline metrics in terms of privacy preservation and naturalness, the advantage it presents in relevance and specificity is less pronounced. Therefore, the development of innovative metrics tailored to specific evaluation criteria presents a valuable avenue for our future research endeavors.
Appendix A appendix
A.1 Question design for human evaluation
Q1: The rewrite deletes/obfuscates _ ?
  • (a)
    None of the key information in the personal information and the original utterance does contain personal information.
  • (b)
    None of the key information in the personal information, because the original utterance does not contain personal information.
  • (c)
    At least one key information in the personal information (if the rewrite uses both correct and incorrect strategies, only evaluate the part that uses the correct strategy).
  • (d)
    All key information in the personal information (using the correct strategy only).
  • (e)
    At least one or all key information in the personal information (using the incorrect strategies only).
Q2: The rewrite _.
  • (a)
    Accurately preserves the meaning of the original sentence.
  • (b)
    Basically the same meaning but does not cover some minor content.
  • (c)
    Has a minor resemblance to the meaning of the original sentence, however, it is also misleading.
  • (d)
    Empty sentence or does not reflect the meaning of the original sentence at all.
Q3: The rewrite is able to retain _ in the original utterance that is not covered in the personal information.
  • (a)
    has no grammatical mistakes and the sentence is coherent.
  • (b)
    has some grammatical mistakes and the sentence is less coherent
  • (c)
    is full of grammatical mistakes</b> and the sentence is not coherent
A.2 Prompt template for synethetic Data
The prompt template used across the paper is shown as 2. We use three nearest examples drawn from the training set as prompting example. Each example contains two cases if the raw persona information is provided. And objective for the prompt is to rewrite given sentence with specified strategy.
Figure 2: Prompt template for T5-NaP2.
A.3 Implementation Details
In our experiment, we consider T5-Base as our targeted rewrite model, we set optimal hyperparameters for model fine tuning with learning rate of $5e^{-4}$ and beam search as decoding method with generative temperature of $0.2$. In the model finetuning, we set noise multiplier of DP-SGD Abadi et al. (2016b) to 0.001 to gain minimal influence for model result. In baseline experiments, for two DP methods applied to echo language model, we consider the empirically optimal noise multipliers 0.01 and epsilon to 3 with one word masked for DPNR. As for DP-Forward-utility, we set the key noise hyperparameters delta to $1e^{-5}$ and epsilon at $7$ to obtain the impact with small noise gap, while for DP-Forward-privacy, we set the hyperparameters to $2e^{-5}$ and $8$ for delta and epsilon respectively. The remaining hyperparemeters are the same as with the ones reported in the corresponding papers.
Appendix B Experiments
B.1 Evaluation metrics
Details of the evaluation metrics for semantic relevance are provided below.
  • MAUVE Pillutla et al. (2021): It is widely used as one of the best automatic metrics to quantify the semantic differences between a pair of texts. We adopt it to assess the distribution divergence between model outputs and human written rewrites.
  • ROUGE-1 Lin (2004): It is a widely used evaluation metric measuring the overlap of unigrams between a generated text and a set of references.
  • ROUGE-Lsum: It is a variant of ROUGE-L, tailored to evaluate longer texts by summarizing the longest common sub-sequences between an output text and a set of references.
B.2 Impact of DP-SGD.
Table 7 shows results of models trained with and without DP-SGD. The purpose is to understand to what degree the widely used DP method can influence rewriting quality if the training data is sensitive. Comparing these two settings with human rewrites, there is a slight performance drop of around 3% with DP-SGD. However, DP-SGD provides a privacy guarantee during training which is useful when the training data is sensitive. When comparing with automatic metrics, as shown in Table 8, there is only a 1% performance drop in terms of privacy leakage if DP-SGD is applied. For preservation of semantic contents, MAUVE scores show little differences between using and not using DP-SGD, meaning our proposed rewriting approaches are compatible with the DP based training algorithms for more sensitive scenarios.
SPrivacy SRel SNatural
Human_deleting 82.00% 76.00% 95.00%
DP_deleting 59.00% 88.00% 99.00%
non-DP_deleting 63.00% 82.00% 96.00%
Human_obscuring 81.00% 97.00% 98.00%
DP_obscuring 29.00% 90.00% 98.00%
non-DP_obscuring 32.00% 88.00% 93.00%
Table 7: Human evaluation results with and without DP-SGD.
DP real synth LLM Privacy_NLI MAUVE ROUGE-1 ROUGE-Lsum
False 1300 0 - 0.0810 $\pm$ 0.1077 0.9468 0.6946 0.6924
False 1300 3900 GPT-3 0.0826 $\pm$ 0.0903 0.9558 0.7143 0.7122
False 1300 3900 GPT-4 0.0619$\pm$0.0870 0.9658 0.7301 0.7278
True 1300 0 - 0.0602 $\pm$ 0.0759 0.9584 0.7338 0.7316
True 1300 3900 GPT-3 0.0757 $\pm$ 0.0908 0.9589 0.7368 0.7351
True 1300 3900 GPT-4 0.0703 $\pm$ 0.1135 0.9514 0.7446 0.7428
Table 8: Evaluation for DP and combination of synthetic data and human rewrites